Let f:R-R be a function such that f (x+y)+f (x-y)=f (xy) for all x,y €R.Then f is:

Grade:12

2 Answers

Arun
4 years ago
 
Put y=0

=> f(x)+f(x)=2f(x).f(0)

=> f(0)=1

Now put x=0 and y=x

=> f(x)+f(-x)=2f(0)f(x)

=> f(x)+f(-x)=2f(x)

=> f(-x)=f(x)

So, even function.
Aditya Gupta
4 years ago
aruns ans is wrong.
we are given that f(x+y) + f(x – y)= f(xy).... (1)
put y=0
then f(x)+f(x)= f(0)
or f(x)= f(0)/2= C (a constant)
so from (1),
C+C=C
or C=0
so that f(x)= 0 is the only possibility.